# Guided Diffusion Sampling on Function Spaces with Applications to PDEs

### v1

@@ B.3 Special Case: Euclidean Setting
This expression is the finite-dimensional analogue of the one we obtain in Eq. 48. This yields
@@
Appendix C Inverse PDE Solver with an FDM Forward Operator
@@
and we formulate the inverse problem as estimating initial state ${\boldsymbol{a}}$ from corrupted observations of solution state ${\boldsymbol{u}}$. Our goal is to sample from $p({\boldsymbol{a}}|{\boldsymbol{u}})$ by applying the function-space reverse diffusion steps with a guidance term $\nabla_{\boldsymbol{a}}\|{\boldsymbol{A}}({\boldsymbol{a}})-{\boldsymbol{u}}\|$. Since ${\boldsymbol{A}}$ is a standard FDM solver, we can compute gradients via automatic differentiation. We present qualitative examples in Figure 5 and show in Table 2 that our method can achieve small relative $L^{2}$-errors in these challenging cases with both FDM as forward operator and joint learning. In practice, considering similar or better performance, we use a joint learning approach for all our experiments as the inference speed is considerably faster, since the FDM-based method requires backpropagating through the linear solve FDM_Solve.
@@
Appendix D Detailed Dataset Description
@@
$-\nabla\cdot({\boldsymbol{a}}(x)\nabla{\boldsymbol{u}}(x))=f(x),\qquad x\in(0,1)^{2},\$ (65)
with constant forcing $f(x)=1$ and zero boundary conditions. We follow the strategies in Li et al. [21] to generate coefficient functions ${\boldsymbol{a}}\sim h_{{}_{\#}}\mathcal{N}(0,(-\Delta+9\mathbf{I})^{-2})$, where $h\colon\mathbb{R}\to\mathbb{R}$ is set to be $12$ for positive numbers and $3$ otherwise.
@@
$\nabla^{2}{\boldsymbol{u}}(x)={\boldsymbol{a}}(x),\quad x\in(0,1)^{2},$ (66)
with homogeneous Dirichlet boundary conditions ${\boldsymbol{u}}|_{\partial\Omega}=0$. We generate piecewise-constant coefficient fields ${\boldsymbol{a}}(x)$ by thresholding Gaussian random fields $\mathcal{N}(0,(-\Delta+9\mathbf{I})^{-2})$ at zero, creating binary patterns of high/low conductivity regions. The PDE guidance function is $f=\nabla^{2}{\boldsymbol{u}}-{\boldsymbol{a}}$.
@@
$\nabla^{2}{\boldsymbol{u}}(x)+k^{2}{\boldsymbol{u}}(x)={\boldsymbol{a}}(x),\quad x\in(0,1)^{2},$ (67)
with $k=1$ and ${\boldsymbol{u}}|_{\partial\Omega}=0$. Coefficient fields ${\boldsymbol{a}}(x)$ are generated as ${\boldsymbol{a}}\sim\mathcal{N}(0,(-\Delta+9\mathbf{I})^{2})$, followed by thresholding to create piecewise constant function. The PDE guidance function is $f=\nabla^{2}{\boldsymbol{u}}+k^{2}{\boldsymbol{u}}-{\boldsymbol{a}}$.
@@
We further evaluate the performance on Navier-Stokes equations by generating its initial and terminal states as in [21]. In particular, we consider the evolution of a vorticity field ${\boldsymbol{u}}(x,t)$ over time given by
$\partial_{t}{\boldsymbol{u}}(x,t)+{\boldsymbol{w}}(x,t)\cdot\nabla{\boldsymbol{u}}(x,t)$ $=\nu\Delta{\boldsymbol{u}}(x,t)+f(x),\quad x\in(0,1)^{2},\,t\in(0,T],$ (68)
$\nabla\cdot{\boldsymbol{w}}(x,t)$ $=0,\quad x\in(0,1)^{2},\,t\in[0,T],$ (69)
@@
where ${\boldsymbol{w}}$ is the velocity field; $\nu=\frac{1}{1000}$, viscosity; and $f$, a fixed forcing term. The initial condition ${\boldsymbol{a}}(x)$ is sampled from $\mathcal{N}(0,7^{3/2}(-\Delta+49\mathbf{I})^{-5/2})$ under periodic boundary conditions. The forcing term is defined as $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$. We simulate the PDE for $T=1$ second with $10$ timesteps using a pseudo-spectral method. As for PDE guidance, we utilize that $\nabla\cdot{\boldsymbol{u}}(x,t)=\nabla\cdot(\nabla\times v)=0$.
@@
We study bounded flow around cylindrical obstacles:
@@
Appendix E Detailed Experiment Setup
@@
We validate our approach by solving both forward and inverse problems on five different PDE problems. These PDEs include Darcy Flow, Poisson, Helmholtz, and Navier-Stokes with and without boundary conditions. We follow the same strategy as in DiffusionPDE [17] to generate datasets, where we prepare $50,000$ training samples and $1,000$ test samples for each PDE. The Navier-Stokes equation with boundary conditions specifically consists of $14,000$ train and $1,000$ test samples. For quantitative comparisons, error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, except for the inverse Darcy Flow problem, where we use binary error rate.
@@
We adopt a 4-level U-shaped neural operator architecture [52] as the denoiser ${\boldsymbol{D}}_{\theta}$, which has 54M parameters, similar to DiffusionPDE’s network size. The network is trained using $50,000$ training samples for 200 epochs. For the multi-resolution training, we begin training on a coarser grid ($64\times 64$) for 200 epochs, then switch to a higher resolution ($128\times 128$) for 100 epochs. The hyperparameters we used for training and inference are listed in Table 3. We source the quantitative results of deterministic baselines from DiffusionPDE [17]’s table.
@@
Appendix F Implementation Details of the Guidance Mechanism
For inference, we tuned the guidance strength $\zeta$ on a small validation set, resulting in the values shown in Table 4. We noted that PDE loss calculations are unreliable in early sampling stages due to high noise levels. Hence empirically, we only apply PDE loss when $\sigma_{t}<1$. Furthermore, to ensure smooth convergence to the posterior, we found it beneficial to dial down the guidance weights as the noise level decreases. Therefore, we implement a simple but effective scheduling scheme for the guidance weights of both observation and PDE loss:
@@ Appendix F Implementation Details of the Guidance Mechanism
We use the Huber loss for PDE guidance instead of mean squared error because it provides robustness against potential outliers caused by finite difference approximation errors, which improves the stability of gradient updates.
@@
PDE Loss Weight 0 0 0 10 1 1 1 10 100 15
Appendix G Additional Experiments
@@
G.1 Plug-and-play inverse solvers
In order to show the adaptability of the framework, we further test our method with various guidance methods. Namely, Table 5 and Figure 6 demonstrate the reconstruction results on Poisson PDE equation with various inverse solvers [40, 39, 63] on different priors. FunDPS corresponds to the combination of Function Space prior with DPS solver and consistently outperforms other methods within a smaller number of steps.
@@ G.1 Plug-and-play inverse solvers
Table 5: The relative errors for Poisson equation under varying priors and plug-and-play inverse solvers. Note that FunDPS corresponds to the intersection of Function Space prior with DPS solver, which outperforms all other methods across different sparse reconstruction tasks within a smaller number of steps. FunDPS results are based on 500 steps, whereas other methods are performed with at least 2000 steps.
Figure 6: The qualitative results of Table 5 with Poisson equation. First and second rows correspond to the forward problem reconstructions and error maps, respectively. The third and fourth rows correspond to the inverse problem reconstructions and error maps, respectively. Relative errors are also reported for this specific data under each error map, where FunDPS achieves the minimum among all the tasks.
G.2 Fully-observed problems
Deep learning approaches have been extensively researched for solving classical forward and inverse problems, where either the initial state or the final state is fully known. While the deterministic baselines are powerful in these tasks, our method still demonstrates superior performance in 3 out of 4 cases, highlighting the strong modeling capabilities of our framework. Ours also outperforms DiffusionPDE in all cases, shown in Table 6.
@@ G.2 Fully-observed problems
DiffusionPDE [^4] We again found reproducibility issues with DiffusionPDE. Our reproduced results are here. Please refer to Appendix I for details. 2000 2.9% 13.0% 2.4% 8.4%
@@
G.3 Mixed-resolution training
Our mixed-resolution training combines efficient low-resolution learning with high-resolution finetuning. We first train for 200 epochs on 64×64 resolution data to learn coarse features, then train for 100 epochs on 128×128 resolution to capture fine details. This approach leverages the resolution independence of neural operators while reducing computational costs.
Table 7 compares training configurations across resolutions. Direct high-resolution training performs well but requires substantially more parameters and computing resources. Our mixed-resolution approach achieves comparable or better performance while maintaining the smaller model size, reducing total GPU hours by about 25%.
@@ G.3 Mixed-resolution training
G.4 Mixed-resolution inference
Thanks to its multi-resolution nature, neural operators can be trained on multiple resolutions and applied to data of different resolutions. We can reduce significant inference time by performing most sampling steps at low resolution and upsampling only near the end to finalize high-frequency details. We found that upsampling in the middle of the diffusion process worked poorly, which we attribute to the difficulty of preserving GRF properties during upscaling. Hence, we propose a bi-level sampling process, ReNoise. The first stage is a complete diffusion sampling at low resolution, from $\sigma^{(1)}_{\max}$ to $\sigma^{(1)}_{\min}$, over $t_{\text{up}}$ steps. Samples are then upscaled (e.g., using bicubic interpolation, though our method is not sensitive to this choice). To counter upscaling artifacts and initiate a subsequent high-resolution diffusion, noise is added at $\sigma=\sigma^{(2)}_{\max}$. Subsequently, a second diffusion sampling process is performed. Empirically, we found that $\sigma^{(2)}_{\min}$ in the initial low-resolution stage need not be very low, as details are refined in the second stage. Furthermore, $\sigma^{(2)}_{\max}$ for initiating the high-resolution process is typically set to a modest value (e.g., in the 1–10 range), considerably lower than the initial $\sigma^{(1)}_{\max}=80$, as we want to keep the existing information in the low-resolution sample.
We provide the multi-resolution inference results in Figure 7. We maintain the same level of accuracy even when up to 80% of the steps run in low resolution, which shows the effectiveness of ReNoise.
G.5 Number of observations
We investigate how the number of observations affects model performance in both forward and inverse Darcy Flow problems. We use 500 steps for evaluation. Figure 8 shows that performance improves consistently as we increase the number of observed points from 100 (0.6%) to 2000 (12%) of the spatial domain. It is worth noting that our method can achieve reasonable accuracy even with extremely sparse observations, demonstrating its effectiveness.
@@ G.5 Number of observations
G.6 Additional experiments on Navier-Stokes with boundary conditions
In Table 1, we present experiments for the Navier–Stokes equations with boundary conditions by using both boundary observations and $1\%$ random interior points. We also compare our methodology against a standard setup in which only $3\%$ of the data is revealed for state reconstruction. Numerical results for this comparison can be found in Table 8.

### v2

@@ C.3 Special Case: Euclidean Setting
This expression is the finite-dimensional analogue of the one we obtain in Eq.˜48. This yields
@@
Appendix D Inverse PDE Solver with an FDM Forward Operator
@@
and we formulate the inverse problem as estimating initial state $\boldsymbol{a}$} from corrupted observations of solution state $\boldsymbol{u}$}. Our goal is to sample from $p({\boldsymbol{a}}|{\boldsymbol{u}})$ by applying the function-space reverse diffusion steps with a guidance term $\nabla_{\boldsymbol{a}}\|{\boldsymbol{A}}({\boldsymbol{a}})-{\boldsymbol{u}}\|$. Since $\boldsymbol{A}$} is a standard FDM solver, we can compute gradients via automatic differentiation. We present qualitative examples in Figure˜5 and show in Table˜2 that our method can achieve small relative $L^{2}$-errors in these challenging cases with both FDM as forward operator and joint learning. In practice, considering similar or better performance, we use a joint learning approach for all our experiments as the inference speed is considerably faster, since the FDM-based method requires backpropagating through the linear solve FDM_Solve.
@@
Appendix E Detailed Dataset Description
@@
$\nabla\cdot(a(x)\nabla u(x))=f(x),\qquad x\in(0,1)^{2},\\$ (64)
with constant forcing $f(x)=1$ and zero boundary conditions. We follow the strategies in Li et al. [21] to generate coefficient functions $a\sim h_{{}_{\#}}\mathcal{N}(0,(-\Delta+9\mathbf{I})^{-2})$, where $h\colon\mathbb{R}\to\mathbb{R}$ is set to be $12$ for positive numbers and $3$ otherwise.
@@
$\nabla^{2}u(x)=a(x),\quad x\in(0,1)^{2},$ (65)
with homogeneous Dirichlet boundary conditions $u|_{\partial\Omega}=0$. We generate piecewise-constant coefficient fields $a(x)$ by thresholding Gaussian random fields $\mathcal{N}(0,(-\Delta+9\mathbf{I})^{-2})$ at zero, creating binary patterns of high/low conductivity regions. The PDE guidance function is $f=\nabla^{2}u-a$.
@@
$\nabla^{2}u(x)+k^{2}u(x)=a(x),\quad x\in(0,1)^{2},$ (66)
with $k=1$ and $u|_{\partial\Omega}=0$. Coefficient fields $a(x)$ are generated as in [17]. The PDE guidance function is $f=\nabla^{2}u+k^{2}u-a$.
@@
We further evaluate the performance on the Navier-Stokes equations by generating its initial and terminal states as in [21]. In particular, we consider the evolution of a velocity field ${\boldsymbol{u}}(x,t)$ over time given by
$\partial_{t}{\boldsymbol{w}}(x,t)+{\boldsymbol{u}}(x,t)\cdot\nabla{\boldsymbol{w}}(x,t)$ $=\nu\Delta{\boldsymbol{w}}(x,t)+f(x),\quad x\in(0,1)^{2},\,t\in(0,T],$ (67)
$\nabla\cdot{\boldsymbol{u}}(x,t)$ $=0,\quad x\in(0,1)^{2},\,t\in[0,T],$ (68)
@@
where ${\boldsymbol{w}}=\nabla\times{\boldsymbol{u}}$ is the vorticity; $\nu=\frac{1}{1000}$, viscosity; and $f$, a fixed forcing term. The initial condition ${\boldsymbol{a}}(x)$ is sampled from $\mathcal{N}(0,7^{3/2}(-\Delta+49\mathbf{I})^{-5/2})$ under periodic boundary conditions. The forcing term is defined as $f(x)=\frac{1}{10}\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$. We simulate the PDE for $T=1$ second using a pseudo-spectral method. As for PDE guidance, we utilize that $\nabla\cdot{\boldsymbol{w}}=\nabla\cdot(\nabla\times{\boldsymbol{u}})=0$.
@@
We study bounded flow around cylindrical obstacles, governed by:
@@
Appendix F Detailed Experiment Setup
@@
We validate our approach by solving both forward and inverse problems on five different PDE problems. These PDEs include Darcy Flow, Poisson, Helmholtz, and Navier-Stokes with and without boundary conditions. We follow the same strategy as in DiffusionPDE [17] to generate datasets, where we prepare $50,000$ training samples and $1,000$ test samples for each PDE. The Navier-Stokes equation with boundary conditions specifically consists of $14,000$ train and $1,000$ test samples. The resolution is $128\times 128$, and in some settings we downsample the data by $2\times$. For quantitative comparisons, error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, except for the inverse Darcy Flow problem, where we use the binary error rate.
@@
We adopt a 4-level U-shaped neural operator architecture [52] as the denoiser ${\boldsymbol{D}}_{\theta}$, which has 54M parameters, similar to DiffusionPDE’s network size. The network is trained using $50,000$ training samples for 200 epochs. For the multi-resolution training, we begin training on a coarser grid ($64\times 64$) for 200 epochs, then switch to a higher resolution ($128\times 128$) for 100 epochs. The hyperparameters we used for training and inference are listed in Table˜3. We source the quantitative results of deterministic baselines from DiffusionPDE [17]’s table.
@@
Appendix G Implementation Details of the Guidance Mechanism
For inference, we tuned the guidance strength $\zeta$ on a small validation set, resulting in the values shown in Table˜4. We noted that PDE loss calculations are unreliable in early sampling stages due to high noise levels. Hence empirically, we only apply PDE loss when $\sigma_{t}<1$. Furthermore, to ensure smooth convergence to the posterior, we found it beneficial to dial down the guidance weights as the noise level decreases. Therefore, we implement a simple but effective scheduling scheme for the guidance weights of both observation and PDE loss:
@@ Appendix G Implementation Details of the Guidance Mechanism
We use the Huber loss for PDE guidance instead of mean squared error because it provides robustness against potential outliers caused by finite difference approximation errors, which improves the stability of gradient updates. We investigated the method’s sensitivity to the guidance strength in Section˜H.6.
@@
PDE Loss Weight 0 0 0 0 1 1 1 10 100 15
Appendix H Additional Experiments
@@
H.1 Plug-and-play inverse solvers
In order to show the adaptability of the framework, we further test our method with various guidance methods. Namely, Table˜5 and Figure˜6 demonstrate the reconstruction results on Poisson PDE equation with various inverse solvers [40, 39, 65] on different priors. FunDPS (Function Space + DPS) consistently outperforms other methods within a smaller number of steps.
The significant underperformance of DDNM and DAPS in our setting is primarily due to the extreme sparsity of observation data. DDNM relies on projecting the sample onto measurement subspace at each step. With only 3% of observation points, the measurement subspace is extremely low-dimensional compared to the overall function space. As a result, the projection provides very weak guidance and leads to poor reconstruction. The core issue of DAPS comes from the localness of guided updates during Langevin dynamics. Only the points with observation are updated and others are just added with noise. This makes the intermediate state after Langevin dynamics discontinuous and out-of-distribution, which hinders performance. If we increase the number of sampling steps to 20,000, the accuracy can match DPS, but with 100x more time.
Table 5: The relative errors for Poisson equation under varying priors and plug-and-play inverse solvers. FunDPS corresponds to the intersection of Function Space prior with DPS solver. FunDPS results are based on 500 steps, whereas other methods are performed with at least 2000 steps.
@@ H.1 Plug-and-play inverse solvers
Figure 6: The qualitative results of Table˜5 with Poisson equation. First and second rows correspond to the forward problem reconstructions and error maps, respectively. The third and fourth rows correspond to the inverse problem reconstructions and error maps, respectively. Relative errors are also reported for this specific data under each error map, where FunDPS achieves the minimum among all the tasks.
H.2 Fully-observed problems
Deep learning approaches have been extensively researched for solving classical forward and inverse problems, where either the initial state or the final state is fully known. While the deterministic baselines are powerful in these tasks, our method still demonstrates superior performance in 3 out of 4 cases, highlighting the strong modeling capabilities of our framework. Ours also outperforms DiffusionPDE in all cases, shown in Table˜6.
@@ H.2 Fully-observed problems
DiffusionPDE [^4] We again found reproducibility issues with DiffusionPDE. Our reproduced results are here. Please refer to Appendix˜I for details. 2000 2.9% 13.0% 2.4% 8.4%
@@
H.3 Multi-resolution training
Our multi-resolution training combines efficient low-resolution learning with high-resolution finetuning. We first train for 200 epochs on 64×64 resolution data to learn coarse features, then train for 100 epochs on 128×128 resolution to capture fine details. This approach leverages the resolution independence of neural operators while reducing computational costs.
Table˜7 compares training configurations across resolutions. Direct high-resolution training performs well but requires substantially more parameters and computing resources. Our multi-resolution approach achieves comparable or better performance while maintaining the smaller model size, reducing total GPU hours by about 25%.
@@ H.3 Multi-resolution training
H.4 Multi-resolution inference
Our multi-resolution sampling pipeline is shown in Figure˜3b. It includes two stages with noise increase in between. The first stage is a complete diffusion sampling at low resolution, from $\sigma^{(1)}_{\max}$ to $\sigma^{(1)}_{\min}$, over $t_{\text{up}}$ steps. Samples are then upscaled (e.g., using bicubic interpolation, though our method is not sensitive to this choice). To counter upscaling artifacts and initiate a subsequent high-resolution diffusion, noise is added at $\sigma=\sigma^{(2)}_{\max}$. Subsequently, a second diffusion sampling process is performed. Empirically, we found that $\sigma^{(2)}_{\min}$ in the initial low-resolution stage need not be very low, as details are refined in the second stage. Furthermore, $\sigma^{(2)}_{\max}$ for initiating the high-resolution process is typically set to a modest value (e.g., in the 1–10 range), considerably lower than the initial $\sigma^{(1)}_{\max}=80$, as we want to keep the existing information in the low-resolution sample. We provide the multi-resolution inference results in Figure˜4. Table˜8 further shows that FunDPS generalizes across resolutions, while DiffusionPDE fails to transfer and must be retrained for each fixed grid.
Table 8: Results on Darcy Flow under different training–inference resolution pairs. DiffPDE-2000 and DiffPDE-500 represent 2000 and 500 steps for Diffusion PDE, respectively. FunDPS, using 500 steps, achieves superior generalization and maintains accuracy under multi-resolution settings, whereas DiffusionPDE fails to generalize between grids.
Training Inference DiffPDE-2000 DiffPDE-500 FunDPS
64 64 6.38% 7.96% 3.03%
64 128 33.72% 35.57% 3.64%
128 128 6.07% 4.60% 2.74%
Mixed 128 3.83% 4.53% 2.49%
H.5 Number of observations
We investigate how the number of observations affects model performance in both forward and inverse Darcy Flow problems. We use 500 steps for evaluation. Figure˜7 shows that performance improves consistently as we increase the number of observed points from 100 (0.6%) to 2000 (12%) of the spatial domain. It is worth noting that our method can achieve reasonable accuracy even with extremely sparse observations, demonstrating its effectiveness.
We further analyze the performance of FunDPS and DiffusionPDE under different observation sparsities on both forward and inverse Darcy Flow problems. As shown in Table˜9, FunDPS maintains high accuracy even with extremely sparse observations (as low as $0.5\%$ of spatial points), while DiffusionPDE performance degrades significantly under sparse settings.
Table 9: Comparison of FunDPS and DiffusionPDE across varying numbers of observations on Darcy Flow. DiffPDE-2000 and DiffPDE-500 represent 2000 and 500 steps for Diffusion PDE, respectively. FunDPS, using 500 steps, maintains <<10% relative error even with as few as 0.5% observed points.
Forward Inverse
# of Obs. DiffPDE-2000 DiffPDE-500 FunDPS DiffPDE-2000 DiffPDE-500 FunDPS
100 7.2% 13.8% 7.2% 10.7% 16.4% 8.0%
200 6.5% 7.8% 4.8% 9.1% 10.2% 6.3%
500 6.1% 4.6% 2.9% 7.9% 8.1% 5.2%
2000 2.2% 2.3% 1.7% 3.9% 4.7% 4.1%
@@ H.5 Number of observations
H.6 Sensitivity to the guidance strength
The guidance weight $\zeta$ is a key hyperparameter in our framework, manually chosen for each task by tuning on a small validation set. This is a common practice in guided diffusion models, as the optimal weight often depends on the forward operator and prior data distribution. We provide a list of tuned $\zeta$ values in Table˜4. In our experience, the model’s performance is stable across wide ranges of $\zeta$. However, values that are too high can cause sampling to become unstable and diverge, while values that are too low result in weak guidance and less accurate reconstructions.
We conducted an ablation study to evaluate performance on Darcy Flow and Poisson equation forward/inverse problems across a range of $\zeta$ values, using 500 sampling steps. The results, shown in Table˜10, demonstrate expected behavior.
Table 10: Ablation study on guidance weight $\zeta$ for Darcy Flow and Poisson problems. Errors are $L^{2}$ relative errors (%). NA indicates divergence due to instability.
$\zeta$ Darcy Forward Darcy Inverse Poisson Forward Poisson Inverse
1000 6.92% 24.02% 3.31% 34.38%
5000 2.92% 11.01% 2.56% 28.05%
10000 2.49% 9.16% 1.99% 19.47%
20000 NA 9.86% 18.86% 20.47%
50000 NA 5.18% NA NA
H.7 Additional experiments on Navier-Stokes with boundary conditions
In Table˜1, we present experiments for the Navier–Stokes equations with boundary conditions by using both boundary observations and $1\%$ random interior points. We also compare our methodology against a standard setup in which only $3\%$ of the data is revealed for state reconstruction. Numerical results for this comparison can be found in Table˜11.